Kendra has $3.15 in dimes and nickels. If she has 4 times as many dimes as nickles, how many of each coin does she have?
Answer by addingup(3677)

dimes: d
nickels: n
d = 4n
0.10d+0.05n = 3.15
0.10(4n)+0.05n = 3.15
0.40n+0.05n = 3.15
0.45n = 3.15
n = 7 she has 7 nickels and 7*4 = 28 dimes
------------------------
check:
28*0.10 = 2.80
7*0.05 = 0.35
_______________
Total . = 3.15 Correct